OpenStudy (anonymous):

find dy/dx at point (1,2) of the implicit function (x^2+y^2)^2=25/4xy^2.... my answer is -11/14.. am i right ? :))

OpenStudy (anonymous):

Show your work.

OpenStudy (anonymous):

\[2(x^2+y^2)(2x+2yy')=\frac{ 25 }{ 4 }.y^2+2yy'.\frac{ 25 }{ 4 }x\]

OpenStudy (anonymous):

then i subtitute (1,2) to (x,y)

OpenStudy (anonymous):

That would be correct then.

OpenStudy (anonymous):

can u pls check it..................... :)

OpenStudy (anonymous):

Ummm, just show the work... I won't do it but I'll check it.

OpenStudy (anonymous):

\[2(1)^2+2(2)^2(2(1)+2(2).y'=\frac{ 25 }{ 4 }(2)+2(2).y'.\frac{ 25 }{ 4 }(1)\] \[2+8(2+4y')=\frac{ 25 }{ 2 }+25y'\] \[2+16+32y'=\frac{ 25 }{ 2 }+25y'\] \[32y'-25y'=\frac{ 25 }{ 2 }-18\] y'=\[y'=-\frac{ 11 }{1 4 }\]

OpenStudy (anonymous):

well , what about this ?

OpenStudy (anonymous):

Wow in the first step you get rid of important parenthesis.

OpenStudy (anonymous):

huh ????? what ?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

\[\begin{array}{rcl} 2(x^2+y^2)(2x+2yy')&=&\frac{ 25 }{ 4 }\cdot y^2+2yy'\cdot\frac{ 25 }{ 4 }x\\ 2((1)^2+(2)^2)(2(1)+2(2)y')&=&\frac{ 25 }{ 4 }\cdot (2)^2+2(2)y'\cdot \frac{ 25 }{ 4 }(1)\\ 2(1+4)(2+4y')&=&25 +y'\cdot 25\\ 2(5)(2+4y')&=&25 +25y'\\ 20+40y'&=&25 +25y'\\ 15y'&=& 5\\ y'&=&\frac{1}{3} \end{array} \]How is this?

OpenStudy (anonymous):

yeahh i miss that part , then the whole thing went worst..

OpenStudy (anonymous):

and wrong...

OpenStudy (anonymous):

@wio thanks a lot !!!!!!!!!! really appreciate it !!!!!!!! ^_^ i got it now :))

OpenStudy (anonymous):

not best response?

OpenStudy (anonymous):

of course best response ! :)